# Development of large-volume 130TeO2 bolometers for the CROSS $2\beta$ decay search experiment

### v1

@@ 3.2 Underground test of large-size 130Te-enriched crystals
To test the newly developed 130Te-enriched crystals as bolometers, we took four of the six samples with a size of 45$\times$45$\times$45 mm each, produced for the CROSS project. Two chosen samples (TeO-20 and TeO-24) were used uncoated, while the other two crystals were metal-coated as follows: i) four-side coated with Al layer (TeO-Al); ii) one-side coated with an AlPd bi-layer in the form of grid (TeO-AlPd). Each crystal was equipped with an NTD-Ge thermistor and a Si:P heater, coupled with UV-curable (PERMABOND 620) and epoxy (Araldite® Rapid) glue respectively. The samples were assembled in the Cu structure with 3D-printed PLA (polylactic acid) spacers — the holder Slim version designed for the CROSS experiment [44]— fasting two crystals per flour. In addition, a thin Ge bolometer (45$\times$45$\times$0.3 mm) has been mounted on each TeO2 crystal to detect Cherenkov radiation. All Ge slabs were SiO-coated to reduce the reflectivity of the surface. A set of Al concentric electrodes, not optimized for the adopted squared geometry of the light detectors, was deposited following the procedure described in [35]. The purpose of the electrodes is to amplify the thermal signal via the Neganov-Trofimov-Luke effect. In total, the tower was composed of 5 floors; the rest of the array was made of six lithium molybdate based bolometers and the same type of Ge light detectors attached to them [45]. The part of the tower containing the 130TeO2 bolometers and construction elements of the detector structure are depicted in figure 8.
@@
The detector array was installed inside a low-background facility with a pulse-tube-based dilution refrigerator, built in the Canfranc underground laboratory (Spain) for the CROSS experiment [46, 47]. In order to mitigate the vibrational noise induced by the pulse tube cryocooler, and to gain space of the available experimental volume, the set-up was recently upgraded with a pendulum-like detector suspension with three magnetic dampers installed at the 1 K stage [48]. Thanks to the underground location of the facility and to massive external and internal shields, the background level from natural radioactivity is well reduced providing suitable conditions for characterisation of large-volume low-temperature detectors [48, 49]. After reaching a base temperature of around 12 mK, we carried out the characterization of the bolometers at warmer temperatures (17, 22, and 27 mK) regulated on the detector plate. The detector was controlled and read-out using room-temperature electronics [50, 51]. The amplified voltage output of the NTD-Ge thermistors was digitized in continuous streams by a 24-bit ADC with a sampling rate of 2 or 10 kHz; the cut-off frequency of the active Bessel filter was set as 0.3 or 2.5 kHz respectively. The optimum-filter-based data processing was done with the same tool, and in a similar way, as described above (section 3.1).
@@
Table 4: Performance of 130Te-enriched TeO2 bolometers (0.55 kg each) at different temperatures of the detector plate ($T_{plate}$). For each working point of the NTD thermistor (resistance $R_{NTD}$ at a given current $I_{NTD}$), we report the pulse-shape time constants (rise- and decay-time parameters, $\tau_{r}$ and $\tau_{d}$), the detector sensitivity represented by the signal voltage amplitude per unit of the deposited energy ($A_{s}$), and the RMS baseline width (RMSnoise).
@@
Detectors calibration and characterization is done with the help of a 232Th source, inserted inside the lead shielding. Coincidences between 130TeO2 detectors and corresponding Ge bolometers have been established according to time properties (rise time) of their thermal signals. An example of a two-dimensional distribution of the amplitudes of the detected events is illustrated in figure 10. A band of events with a light signal amplitude clearly depending on the heat energy — appreciable in figure 10 — corresponds to $\gamma$($\beta$) interactions in the 130TeO2 absorber resulting to the emission of Cherenkov radiation. Clusters of high-energy particles with no detectable light signals correspond to events induced by $\alpha$ decays; a possible tiny scintillation of the TeO2 material is not observed for $\alpha$-particles as the conditions of light collection are not favorable due to the absence of a reflective film around the crystal. This is also the main reason for only partial discrimination between $\alpha$ and $\gamma$($\beta$) populations. We notice in fact that a much better discrimination was obtained with a device consisting of an even larger (50-mm cube) crystal placed inside a holder with a reflective foil [36], characterized by a similar-performance but smaller light detector. The discrimination will be improved by using an optimized geometry of the Al electrode with a better Ge-wafer surface coverage, which will allow us to improve the amplification by a factor 2 [45]. Moreover, a further enhancement in the NTL gain can be achieved by facing grids to the crystal surface and by applying a twice higher electrode bias [45].
@@
Figure 11: Energy spectra of events detected by the 0.55-kg 130TeO2 bolometers TeO-20 (a), TeO-AlPd (b), and TeO-Al (c). Both metal-coated bolometers were additionally irradiated by a close $\alpha$ source of 238U/234U. The data were acquired in the CROSS set-up at the Canfranc underground laboratory in the following conditions: (a) and (b) correspond to measurements at 27 mK over 116 h (reference date is Nov. 2023); (c) represents 168-h-long dataset collected at 17 mK (reference date is Mar. 2023). The origin of the most intense $\gamma$ and $\alpha$ peaks observed in the spectra is labeled.
@@ Acknowledgments
This work is supported by the European Commission (Project CROSS, Grant No. ERC-2016-ADG, ID 742345), by the Agence Nationale de la Recherche (ANR France; Project CUPID-1, ANR-21-CE31-0014), and US National Science Foundation (NS 1614611). We acknowledge also the support of the P2IO LabEx (ANR-10-LABX0038) in the framework “Investissements d’Avenir” (ANR-11-IDEX-0003-01 – Project “BSM-nu”) managed by ANR, France. This work was also supported by the National Research Foundation of Ukraine under Grant No. 2020.02/0011 and by the National Academy of Sciences of Ukraine in the framework of the project “Development of bolometric experiments for the search for double beta decay”, the grant number 0121U111684. Russian and Ukrainian scientists have given and give crucial contributions to CROSS. For this reason, the CROSS collaboration is particularly sensitive to the current situation in Ukraine. The position of the collaboration leadership on this matter, approved by majority, is expressed at https://a2c.ijclab.in2p3.fr/en/a2c-home-en/assd-home-en/assd-cross/.

### v2

@@ 3.2 Underground test of large-size 130Te-enriched crystals
To test the newly developed 130Te-enriched crystals as bolometers, we took randomly four of the six cubic samples (45 mm side each) produced for the CROSS project. Two chosen samples (TeO-20 and TeO-24) were used uncoated, while the other two crystals were metal-coated as follows: i) four-side coated with a 200 nm Al layer (TeO-Al); ii) one-side coated with an AlPd bi-layer, 100 nm / 10 nm, in the form of a grid (TeO-AlPd). Each crystal was equipped with an NTD-Ge thermistor and a Si:P heater, coupled with UV-curable (PERMABOND 620) and epoxy (Araldite® Rapid) glue respectively. The samples were assembled in the Cu structure with 3D-printed PLA (polylactic acid) spacers — the holder Slim version designed for the CROSS experiment [48]— fasting two crystals per floor. In addition, a thin Ge bolometer (45 $\times$ 45 $\times$ 0.3 mm) has been mounted on each TeO2 crystal to detect Cherenkov radiation. All Ge slabs were SiO-coated (70 nm) to reduce the reflectivity of the surface. A set of Al concentric electrodes, not optimized for the adopted squared geometry of the light detectors, was deposited following the procedure described in [37]. The purpose of the electrodes is to amplify the thermal signal via the Neganov-Trofimov-Luke effect, as used in the above presented measurements (see section 3.1). In total, the tower was composed of 5 floors; the rest of the array was made of six Li2MoO4-based bolometers and the same type of Ge light detectors attached to them [49]. The part of the tower containing the 130TeO2 bolometers and construction elements of the detector structure are depicted in figure 8.
@@
The detector array was installed inside a low-background facility with a pulse-tube-based dilution refrigerator, built in the Canfranc underground laboratory (Spain) for the CROSS experiment [50, 51]. In order to mitigate the vibrational noise induced by the pulse tube cryocooler, and to gain space of the available experimental volume, the set-up was recently upgraded with a pendulum-like detector suspension with three magnetic dampers installed at the 1 K stage [52]. Thanks to the underground location of the facility and to massive external and internal shields, the background level from natural radioactivity is well reduced providing suitable conditions for characterisation of large-volume low-temperature detectors [52, 53]. After reaching a base temperature of around 12 mK, we carried out the characterization of the bolometers at warmer temperatures (17, 22, and 27 mK) regulated on the detector plate. The detector array was controlled and read-out using room-temperature electronics [54, 55]. The amplified voltage output of the NTD-Ge thermistors was digitized in continuous streams by a 24-bit ADC with a sampling rate of 2 or 10 kHz; the cut-off frequency of the active Bessel filter was set as 0.3 or 2.5 kHz respectively. The optimum-filter-based data processing was done with the same tool, and in a similar way, as described above (section 3.1).
@@
Table 4: Performance of 130Te-enriched TeO2 bolometers (0.55 kg each) at different temperatures of the detector plate ($T_{plate}$). For each working point of the NTD thermistor (resistance $R_{NTD}$ at a given current $I_{NTD}$), we report the pulse-shape time constants ($\tau_{r}$ and $\tau_{d}$), the detector sensitivity ($A_{s}$; evaluated with 2615 keV $\gamma$ quanta), and the RMS baseline width (RMSnoise).
@@
Detectors calibration and characterization is done with the help of a 232Th source, inserted inside the lead shielding. Coincidences between 130TeO2 detectors and corresponding Ge bolometers have been established according to time properties (rise time) of their thermal signals. An example of a two-dimensional distribution of the amplitudes of the detected events is illustrated in figure 10. A band of events with a light signal amplitude clearly depending on the heat energy — appreciable in figure 10 — corresponds to $\gamma$($\beta$) interactions in the 130TeO2 absorber resulting to the emission of Cherenkov radiation. Clusters of high-energy particles with no detectable light signals correspond to events induced by $\alpha$ decays; a possible tiny scintillation of the TeO2 material is not observed for $\alpha$-particles as the conditions of light collection are not favorable due to the absence of a reflective film around the crystal. This is also the main reason for only partial discrimination between $\alpha$ and $\gamma$($\beta$) populations [^2] The impact of a single-side grid coating on the photon absorption and thus on photon output from the crystal is expected to be negligibly low, as can be concluded from studies of metal-coated lithium molybdate scintillating bolometers [27, 28]. On the contrary, the 4-side Al coating can lead to the efficient absorption of photons by the film resulting to a poor light output, as observed for lithium molybdate crystals [58].. We notice in fact that a much better discrimination was obtained with a device consisting of an even larger (50-mm cube) crystal placed inside a holder with a reflective foil [38], characterized by a similar-performance but smaller light detector. The discrimination will be improved by using an optimized geometry of the Al electrode with a better Ge-wafer surface coverage, which will allow us to improve the amplification by a factor 2 [49]. Moreover, a further enhancement in the NTL gain can be achieved by facing grids to the crystal surface and by applying a twice higher electrode bias [49].
@@
Figure 11: Energy spectra of events detected by the 0.55-kg 130TeO2 bolometers TeO-20 (a), TeO-AlPd (b), and TeO-Al (c). Both metal-coated bolometers were additionally irradiated by a close $\alpha$ source of 238U/234U. The data were acquired in the CROSS set-up at the Canfranc underground laboratory in the following conditions: (a) and (b) correspond to measurements at 27 mK over 116 h (reference date is Nov. 2023); (c) represents a 168-h-long dataset collected at 17 mK (reference date is Mar. 2023). The origin of the most intense $\gamma$ and $\alpha$ peaks observed in the spectra is labeled.
@@ Acknowledgments
This work is supported by the European Commission (Project CROSS, Grant No. ERC-2016-ADG, ID 742345), by the Agence Nationale de la Recherche (ANR France; Project CUPID-1, ANR-21-CE31-0014), and US National Science Foundation (NS 1614611). We acknowledge also the support of the P2IO LabEx (ANR-10-LABX0038) in the framework “Investissements d’Avenir” (ANR-11-IDEX-0003-01 – Project “BSM-nu”) managed by ANR, France. This work was also supported by the National Research Foundation of Ukraine under Grant No. 2023.03/0213 and by the National Academy of Sciences of Ukraine in the framework of the project “Development of bolometric experiments for the search for double beta decay”, the grant number 0121U111684. Russian and Ukrainian scientists have given and give crucial contributions to CROSS. For this reason, the CROSS collaboration is particularly sensitive to the current situation in Ukraine. The position of the collaboration leadership on this matter, approved by majority, is expressed at https://a2c.ijclab.in2p3.fr/en/a2c-home-en/assd-home-en/assd-cross/.